OpenStudy (anonymous):

Solve for the variable. x + 80 = -19

OpenStudy (btaylor):

subtract 80 from both sides --> x = -19 - 80 = -99

OpenStudy (anonymous):

x=-19-80 x=-99

OpenStudy (anonymous):

thx